# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ Appendix B Proofs
Armed with Assumption 3.3, we first state some useful inequalities that will form the core of the proof for Assumption 3.3
@@
$\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $=\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]$ (18)
$\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$ (19)
$\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbb{E}[Y\mid X=x,D\in\tilde{\mathcal{S}}_{<}(x)]$ (20)
$\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$ (21)
$\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbb{E}[Y\mid X=x,D\in\tilde{\mathcal{S}}_{>}(x)]$ (22)
$\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$ (23)
@@
For Eq. 18, the event $\{D\in\mathcal{S}_{=}(x)\}$ implies $\{A=\pi_{e}(x),M=f_{M}(\pi_{e})\}$ from the definition of $\mathcal{S}_{=}(x)$ (see Assumption 3.3). Moreover, we have it that $Y\perp\!\!\!\perp D\mid X,A,M$ from Assumption 2.1. Hence, $\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]=E[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$, and this conditional expectation is well-defined when $x\in\mathcal{X}_{=}$. The stated equality follows.
For Eq. 19, the event $\{D\in\mathcal{S}_{<}(x)\}$ implies $\{A=\pi_{e}(x),M<f_{M}(\pi_{e})\}$ from the definition of $\mathcal{S}_{<}(x)$ (see Assumption 3.3). By the independence $Y\perp\!\!\!\perp D\mid X,A,M$, we have it that $\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]=E[Y\mid X=x,A=\pi_{e},M<f_{M}(\pi_{e})]$, and this conditional expectation is well-defined when $x\in\mathcal{X}_{<}$. Finally, we make use of Assumption 3.1, which implies that $\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbb{E}[Y\mid X=x,A=\pi_{e},M<f_{M}(\pi_{e})]$. The stated inequality follows.
For Eq. 20, the event $\{D\in\tilde{\mathcal{S}}_{<}(x)\}$ implies $\{A=\pi_{e}(x),M<f_{M}(\pi_{e})\}$ from the definition of $\tilde{\mathcal{S}}_{<}(x)$ (see Assumption 3.3). The rest follows from the same steps as the paragraph above.
For Eqs. 21 and 22, the arguments are analogous to the arguments for Eqs. 19 and 20, where $\mathcal{X}_{<},\mathcal{S}_{<}$ are replaced by $\mathcal{X}_{>},\mathcal{S}_{>}$, and the direction of the inequality flips since, under Assumption 3.1, $\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbb{E}[Y\mid X=x,A=\pi_{e},M>f_{M}(\pi_{e})]$.
@@
$\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x)]$ (24)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x),M]\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x)]$ (25)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),M]\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x)]$ (26)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=a_{0},M]\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x)]$ (27)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=a_{0}]\mid X=x,A=\pi_{e}(x),D\in\mathcal{S}_{*}(x)]$ (28)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0}]$ (29)
$=\mathbf{1}\{x\in\mathcal{X}_{*},\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ (30)
where Eq. 24 follows from the fact that the event $\{D\in\mathcal{S}_{*}(x)\}\implies\{A=\pi_{e}(x)\}$ by the definition of $\mathcal{S}_{*}$ (see Assumption 3.3). Equation 25 simply applies the law of total probability, including $M$ in the inner expectation (and where we use $\mathbb{E}_{M}$ as helpful shorthand to remind the reader that the outer expectation is taken over $M$). Equation 26 uses the fact that $Y\perp\!\!\!\perp D\mid A,M,X$ in our data-generating process to remove $D$ from the inner expectation. Equation 27 replaces $\pi_{e}(x)$ with $a_{0}$ due to the outside indicator that restricts to $x$ where $\pi_{e}(x)=a_{0}$, so whenever this expression is non-zero, then $\pi_{e}(x)=a_{0}$. Finally, Eq. 28 uses Section B.1, which implies that $\mathbb{E}[Y\mid X=x,A=a_{0},M=m]=\mathbb{E}[Y\mid X=x,A=a_{0}]$, and Eq. 29 uses the fact that the inner expectation is a constant value, to remove the outer expectation over $M$. From Eq. 29, we can simply add back the conditioning on $M=f_{M}(\pi_{e})$, again using Section B.1, and recall that $\pi_{e}=a_{0}$ under the indicator, to arrive at Eq. 30, which completes the proof. ∎
@@
$\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}Y_{\text{min}}$
$\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}Y_{\text{max}}$
$\mathbf{1}\{x\not\in\mathcal{X}_{*},\pi_{e}(x)=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{x\not\in\mathcal{X}_{*},\pi_{e}(x)=a_{0}\}Y_{\text{min}}$
$\mathbf{1}\{x\not\in\mathcal{X}_{*},\pi_{e}(x)=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{x\not\in\mathcal{X}_{*},\pi_{e}(x)=a_{0}\}Y_{\text{max}}$
@@
11 11 $=\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}$ (31)
Where we recall that $\mathcal{X}_{=},\mathcal{X}_{<},\mathcal{X}_{\not\leq}$ form a partition over $X$ by construction. Hence, we can write that
@@
$\quad=\mathbb{E}[(\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}$
$\qquad\qquad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\})\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]]$ (34)
$\quad=\mathbb{E}\big{[}\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\big{]}$ (35)
where Eq. 32 follows from the law of iterated expectation, Eq. 33 follows from consistency (Corollary A.1) and the fact that $Y(a,m)\perp\!\!\!\perp A,M\mid X$ (Corollary A.2), and Eq. 34 follows from Eq. 31. After distributing terms in Eq. 35, the lower bound follows from the application of Sections B.1 and B.1 to each term.
First, $\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(x)]$ follows from Section B.1. Note that Section B.1 applies for all values of $A$, so restricting the possible values of $A$ to $A=\pi_{e}(x)\neq a_{0}$ does not affect its correctness.
Next, $\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{<}(x)]$ follows from Section B.1. Alternatively, $\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\tilde{\mathcal{S}}_{<}(x)]$, where we replace $\mathcal{S}_{<}(x)$ with $\tilde{\mathcal{S}_{<}(x)}$, is also true from Section B.1.
Next, $\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}Y_{min}$ follows from Section B.1.
Next, $\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{*}(x)]$ follows from Section B.1.
Finally, $\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}Y_{min}$ follows from Section B.1.

### v2

@@ Appendix D Proofs
Armed with Assumption 3.3, we first state some useful inequalities that will form the core of the proof for Remark 3.1
@@
$\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\neq\varnothing\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\neq\varnothing\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$ (22)
$\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\neq\varnothing\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\neq\varnothing\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$ (23)
$\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $=$
$\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=$ $a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$ (24)
@@
For Eq. 22, the event $\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\}$ implies $\{A=\pi_{e}(x),M\leq f_{M}(\pi_{e})\}$ from the definition of $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)$ (see Assumption 3.3). By the independence $Y\perp\!\!\!\perp\Pi\mid X,A,M$, we have it that $\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]=E[Y\mid X=x,A=\pi_{e},M\leq f_{M}(\pi_{e})]$, and this conditional expectation is well-defined when $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\neq\varnothing$. Finally, we make use of Assumption 3.1, which implies that $\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbb{E}[Y\mid X=x,A=\pi_{e},M\leq f_{M}(\pi_{e})]$. The stated inequality follows.
For Eq. 23, the event $\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\}$ implies $\{A=\pi_{e}(x),M\geq f_{M}(\pi_{e})\}$ from the definition of $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)$ (see Assumption 3.3). By the independence $Y\perp\!\!\!\perp\Pi\mid X,A,M$, we have it that $\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]=E[Y\mid X=x,A=\pi_{e},M\geq f_{M}(\pi_{e})]$, and this conditional expectation is well-defined when $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\neq\varnothing$. Finally, we make use of Assumption 3.1, which implies that $\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbb{E}[Y\mid X=x,A=\pi_{e},M\geq f_{M}(\pi_{e})]$. The stated inequality follows.
@@
$\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x)]$ (25)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x),M]\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x)]$ (26)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=\pi_{e}(x),M]\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x)]$ (27)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=a_{0},M]\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x)]$ (28)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}_{M}[\mathbb{E}[Y\mid X=x,A=a_{0}]\mid X=x,A=\pi_{e}(x),\Pi\in\mathbf{\Pi}^{e}(x)]$ (29)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0}]$ (30)
$=\mathbf{1}\{\mathbf{\Pi}^{e}(x)\neq\varnothing,\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ (31)
where Eq. 25 follows from the fact that the event $\{\Pi\in\mathbf{\Pi}^{e}(x)\}\implies\{A=\pi_{e}(x)\}$ by the definition of $\mathbf{\Pi}^{e}(x)$ (see Assumption 3.3). Equation 26 simply applies the law of total probability, including $M$ in the inner expectation (and where we use $\mathbb{E}_{M}$ as helpful shorthand to remind the reader that the outer expectation is taken over $M$). Equation 27 uses the fact that $Y\perp\!\!\!\perp\Pi\mid A,M,X$ in our data-generating process to remove $\Pi$ from the inner expectation. Equation 28 replaces $\pi_{e}(x)$ with $a_{0}$ due to the outside indicator that restricts to $x$ where $\pi_{e}(x)=a_{0}$, so whenever this expression is non-zero, then $\pi_{e}(x)=a_{0}$. Finally, Eq. 29 uses Section D.1, which implies that $\mathbb{E}[Y\mid X=x,A=a_{0},M=m]=\mathbb{E}[Y\mid X=x,A=a_{0}]$, and Eq. 30 uses the fact that the inner expectation is a constant value, to remove the outer expectation over $M$. From Eq. 30, we can simply add back the conditioning on $M=f_{M}(\pi_{e})$, again using Section D.1, and recall that $\pi_{e}=a_{0}$ under the indicator, to arrive at Eq. 31, which completes the proof. ∎
@@
$\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}Y_{\text{min}}$
$\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbb{E}[Y\mid X=x,A=\pi_{e},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}Y_{\text{max}}$
$\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(x)=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0},M=f_{M}(\pi_{e})]$ $\geq\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(x)=a_{0}\}Y_{\text{min}}$
$\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(x)=a_{0}\}\mathbb{E}[Y\mid X=x,A=a_{0},M=f_{M}(\pi_{e})]$ $\leq\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(x)=a_{0}\}Y_{\text{max}}$
@@
11 11 $=\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}$ (32)
Here, the subsets of $X\in\mathcal{X}$ satisfying each of $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ and $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}$ form a partition over $\mathcal{X}$ by definition as does $\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}$ and $\{\mathbf{\Pi}^{e}(X)=\varnothing\}$. Hence, we can write that
@@
$\quad=\mathbb{E}[(\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}$
$\qquad\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\})\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]]$ (35)
$\quad=\mathbb{E}\big{[}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]$
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\big{]}$ (36)
where Eq. 33 follows from the law of iterated expectation, Eq. 34 follows from consistency (Corollary A.1) and the fact that $Y(a,m)\perp\!\!\!\perp A,M\mid X$ (Corollary A.2), and Eq. 35 follows from Eq. 32. After distributing terms in Eq. 36, the lower bound follows from the application of Sections D.1 and D.1 to each term.
First, $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$ follows from Section D.1.
Next, $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}Y_{min}$ follows from Section D.1.
Next, $\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(x)]$ follows from Section D.1.
Finally, $\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\geq\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}Y_{min}$ follows from Section D.1.